each exterior of a polygon measures 45 degrees find the sum of the measures of the polygons interior angels
Answer by Alan3354(69443)

The sum of exterior angles for ALL polygons is 360 degs.
360/45 = 8 sides
8*(180 - 45) = 1080 degs
--------
Or, 180*(n-2) = 180*6 = 1080 degs